SINGLE-PHOTON VACUUM ULTRAVIOLET RYDBERG TAGGING TIME-OF-FLIGHT STUDY OF NASCENT O(^{3}P_{2,1,0}) FORMED IN THE 193.3 NM PHOTODISSOCIATION OF SO_{2}

\underline{BRANT JONES}, JINGANG ZHOU, LEI YANG, C. Y. NG, Univeristy of California, Davis One Shields Avenue Davis, CA 95616.

Single-photon Rydberg tagging time-of-flight study on oxygen atom has been firstly demonstrated. The pulsed field ionization photoion (PFI-PI) spectrum for oxygen atoms O(^{3}P_{2,1,0}) resulting from the 193.3 nm photodissociation of SO_{2} have been measured using tunable vacuum ultraviolet (VUV) laser radiation in the frequency range of 109 200 - 110 000 cm^{−1}. The PFI-PI measurement reveals over 120 Rydberg lines, which have been assigned as Rydberg states [2s^{2}2p^{3} (^{4}S^{o}_{3/2}) nd ^{3}D (n = 12-62)] converging to the ground ionic state O^{+}(^{4}S^{o}_{3/2}) formed by the VUV excitation of O(^{3}P_{2,1,0}). The identification of these Rydberg series has led to the development of the single-photon O-atom Rydberg tagging time-of-flight method. Oxygen atoms excited to high-n Rydberg levels were field ionized at the detector. As expected, the translational energy distribution obtained from oxygen atom Rydberg tagging time-of-flight measurement has a higher energy resolution than, but agrees with that derived from the velocity mapped images of the SO/O radical photofragments. Both confirm vibrational structure related to the formation of SO in the ν = 0,1,2 levels, with ν = 2 dominating the partition of available energy. This novel single-photon VUV-excited atom Rydberg tagging technique may prove to be universally applicable to other atoms, and thus become a promising avenue for future photodissociation and reaction dynamic studies.